the volume of a gas at temperature -73°c and pressure 480mm hg is 1000 cm find the volume of the same gas at pressure 760mm of maccury and 127°c


1
Expert's answer
2021-02-25T12:28:26-0500

"T_1 = -73 + 273 = 200 \;K \\

P_1 = 480 \;mmHg \\ = 0.631 \;atm

V_1 = 1000 \;cm^3 = 1 \;L \\

T_2 = 127 + 273 = 400 \;K \\

P_2 = 760 \;mmHg = 1 \;atm \\

\frac{P_1 \times V_1}{T_1} = \frac{P_2 \times V_2}{T_2} \\

V_2 = \frac{P_1 \times V_1 \times T_2}{T_1 \times P_2} \\

= \frac{0.631 \times 1 \times 400}{200 \times 1} \\

= 1.262 \;L = 1262 \;cm^3"